What will happen if a magnetic field and a electric field is place 90 degree angle???

Grade:12th pass

1 Answers

Susmita
5 years ago
Let the electric field be along the z direction and B along x direction.
If in such a field a charged particle of charge q enters with a velocity v=vyj+vzk,then force on it will be
F=q(E+v×B)
Where B=Bi and E=Ek.
If you calculate it you will see the particle will exhibit a cycloid ,ie,a circlular path whose centre move with a speed u=E/B.